OpenStudy (kirbykirby):

How to find the period of a Markov chain? I have the following one-step transition probability matrix:\[P= \begin{bmatrix} 1/2 & 1/2 & 0 & 0 & 0 \\[0.3em] 1/2 & 1/2 & 0 & 0 & 0 \\[0.3em] 0 & 0 & 1/2 & 1/2 & 0 \\[0.3em] 0 & 0 & 1/2 & 1/2 & 0 \\[0.3em] 1/4 & 1/4 & 0 & 0 & 1/2 \end{bmatrix}\] FYI, the definition of the period, \(d_i\), is \(d_i={g.c.d.} \left\{n \in \mathbb{Z}^{+} \mid P_{i,i}^{(n)}>0\right\}\), where \(P_{i,i}^{(n)}=P(X_n=j\mid X_0=i)\).

OpenStudy (blockcolder):

Shouldn't that last one be either \(P_{i,j}^{(n)}\) or \(P_{j,i}^{(n)}\)?

OpenStudy (kirbykirby):

no I checked in different sources and it's definitely \(P_{i,i}^{(n)}\)

OpenStudy (blockcolder):

But that's odd. If I interpreted it correctly, then \(P_{i,i}^{(n)}\) is the probability that the value of the random variable \(X_n=j\) given that initially \(X_0=i\). I just thought it would make more sense if that were denoted as \(P_{i,j}^{(n)}\), since \(a_{ij}\) (ith row, jth column of P) is the probability that X moves from state i to state j. To start the solution, consider states 1, 2, and 5 separately from states 3 and 4.

OpenStudy (blockcolder):

Notice that if X is in state 1, 2, or 5, then it won't go to states 3 and 4. Also, if X is in satate 3 or 4, then it will never states 1, 2, and 5. Noticing this, we let \[A=\begin{bmatrix}1/2&1/2&0\\1/2&1/2&0\\1/4&1/4&1/2\end{bmatrix}; B=\begin{bmatrix}1/2&1/2\\1/2&1/2\end{bmatrix}\] Let us start with B, for now. Our goal is to find \(P_{i,i}^{(n)}\) for B. Remember that \(B^n\) is the transition matrix from state i to state j in n steps. Thus, finding \(P_{i,i}^{(n)}\) is equivalent to finding \(a_{ii}\) of \(B^n\).

OpenStudy (blockcolder):

We can use the Cayley-Hamilton Theorem to find a general formula for B^n. The characteristic equation of B is \(p(\lambda)=\det{(\lambda I-B)}=\lambda^2-\lambda\). Thus, it is true that \(p(B)=0\Rightarrow B^2-B=0\Rightarrow B^2=B\). If we multiply both sides of this last equality by B, we get \(B^3=B^2=B\), etc. Thus, \(B^n=B\) and \(P_{i,i}^{(n)}=1/2\) for states 3 and 4.

OpenStudy (kirbykirby):

Oh dear I am so sorry :( I didn't realize my mistake was in the condition probability notation aye... P(Xn = i | Xo = i) :(

OpenStudy (blockcolder):

Anyway, my solution was made with that consideration. So don't worry. :)

OpenStudy (blockcolder):

For the matrix A, the entries \(a_{ii}\) of \(A^n\) can be found by induction. If you look at A^2: \[A^2=\begin{bmatrix}1/2&1/2&0\\1/2&1/2&0\\3/8&3/8&1/4\end{bmatrix}\] then you can see that \(P_{5,5}^{(2)}=1/4\). Continuing, \[A^3=\begin{bmatrix}1/2&1/2&0\\1/2&1/2&0\\7/16&7/16&1/8\end{bmatrix}\] Thus, \(P_{5,5}^{(3)}=1/8\) and generally, \(P_{5,5}^{(n)}=1/2^n>0~\forall n\). Also, \(P_{1,1}^{(n)}=P_{2,2}^{(n)}=1/2>0.\) Now, all you have to do is to find the gcd of all n where the probability of X returning to its initial state is nonzero. Since each \(P_{i,i}^{(n)}>0\) for all i and n, the last question to answer is: What is the gcd of 1, 2, 3, ...?

OpenStudy (kirbykirby):

Oh that was very helpful :D thank you!! That was actually an interesting way of using the Cayley-Hamilton theorem (though honestly I barely have any experience using it..) and the gcd would be 1then

OpenStudy (blockcolder):

It just so happened that the characteristic polynomial of B was really simple. If it were complicated, I would have resorted to finding a pattern and using induction, like I did with A.

OpenStudy (kirbykirby):

True, but it was a nice result :) I don't think I would have ever though of using it, ever.. :P

OpenStudy (kirbykirby):

thought*